OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomials x4 -54x2 + 576 = 0 A. no solution B. -4, -6 C. 0,4,6 D. -4,4,-6,6 It's not A!

OpenStudy (micahm):

yeah

OpenStudy (anonymous):

do you know the answer?

OpenStudy (micahm):

x4 -54x2 + 576 = 0 what does that look like could you draw that from me

OpenStudy (anonymous):

no

OpenStudy (micahm):

what the name of the lesson

OpenStudy (anonymous):

Sorry, I don't come out to any of the answers. I got 1,0

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

Np :) @DJ♬Øϟϟ¥

OpenStudy (micahm):

B or D

OpenStudy (anonymous):

ok

OpenStudy (micahm):

A is wrong

OpenStudy (anonymous):

yes

OpenStudy (micahm):

C looks wrong

OpenStudy (anonymous):

okay so it between b & d

OpenStudy (micahm):

D

OpenStudy (micahm):

x^4-52x^2+576 = (x^2-16)(x^2-36) So, either (x^2-16) = 0, or (x^2-36) = 0. Hence, refactor and you will get (x-4)(x-4) = 0, or (x-6)(x-6) = 0

OpenStudy (micahm):

that's D

OpenStudy (anonymous):

ok thank u